OpenStudy (anonymous):

use logarithmic differentiation to find the derivative of the function. y = (sqrt(xe^x^2))(x^2+1)^7

zepdrix (zepdrix):

\[\large y=\sqrt{x e^{x^2}}(x^2+1)^7\] Is this what our function looks like?

OpenStudy (anonymous):

yea

zepdrix (zepdrix):

Taking the log of both sides,\[\large \ln y=\ln\left[\sqrt{x e^{x^2}}(x^2+1)^7\right]\]We can use rules of logarithms to break apart the right side. Here is one rule we'll be using: \(\large \log(ab)=\log a + \log b\)

zepdrix (zepdrix):

Which allows us to split it up like this: \[\large \ln y=\ln\left[\sqrt{x e^{x^2}}\right]+\ln\left[(x^2+1)^7\right]\]

zepdrix (zepdrix):

This first term we can umm, rewrite the square root as a rational expression:\[\large \ln\left[\sqrt{x e^{x^2}}\right] \qquad=\qquad \ln\left[({x e^{x^2}})^{1/2}\right]\]Applying the exponent to everything inside the brackets:\[\large = \qquad \ln\left[x^{1/2}e^{x^2/2}\right]\]Then we can use our rule from earlier to break this down further:\[\large = \qquad \ln\left[x^{1/2}\right]+\ln\left[e^{x^2/2}\right]\] Understand what's going on here? :o Logarithmic differentitation can be a pain in the butt. Doing this the normal way would've been better. But they told us to use this process :P

OpenStudy (anonymous):

lol. i am stating to understand it better. breaking it down like this is helping alot

zepdrix (zepdrix):

So we're left with this so far:\[\large \ln y= \qquad \ln\left[x^{1/2}\right]+\ln\left[e^{x^2/2}\right]+\ln\left[(x^2+1)^7\right]\] We want to apply another rule of exponents: \(\large \log(a^b)=b\cdot\log(a)\) Applying this to our first term gives us:\[\large \ln\left[x^{1/2}\right] \qquad=\qquad \frac{1}{2}\ln x\]

zepdrix (zepdrix):

Woah I dunno why that first line is bold like that :o lol

OpenStudy (anonymous):

lol no worries

zepdrix (zepdrix):

\[\large \ln\left[(e)^{x^2/2}\right] \qquad=\qquad \frac{x^2}{2}\ln\left[e\right]\] Do you know what the natural log of e simplifies down to? :o

OpenStudy (anonymous):

do u mean the derivative of ln[e]

zepdrix (zepdrix):

No, just the value of ln e.

zepdrix (zepdrix):

If you don't remember, throw it into your calculator :3

OpenStudy (anonymous):

its 1

zepdrix (zepdrix):

Ok cool, so that simplifies our middle term to:\[\large \frac{x^2}{2}\ln\left[e\right] \qquad=\qquad \frac{x^2}{2}\]

OpenStudy (anonymous):

makes sense

zepdrix (zepdrix):

For our last term, we'll apply this second rule of logs again:\[\large \ln\left[(x^2+1)^7\right] \qquad=\qquad 7\ln\left[x^2+1\right]\]

zepdrix (zepdrix):

So this is what we have so far:\[\large \ln y= \qquad \frac{1}{2}\ln x+\frac{x^2}{2}+7\ln\left[x^2+1\right]\] All we've done so far is applied a bunch of rules to make the problem easier to differentiate. Now that it's easier to work with, we'll apply the derivative from here.

OpenStudy (anonymous):

alright

zepdrix (zepdrix):

Are you comfortable with taking the derivative of logs? Do you know what the left side gives us when we differentiate?

OpenStudy (anonymous):

1/y

zepdrix (zepdrix):

Good. Since we're taking the derivative of y, with respect to x, we'll also get a y' when we take the derivative of that term. \[\large \left(\ln y\right)' \qquad=\qquad \frac{1}{y}y'\]

zepdrix (zepdrix):

\[\large \frac{1}{y}y'= \qquad \left(\frac{1}{2}\ln x+\frac{x^2}{2}+7\ln\left[x^2+1\right]\right)'\] How bout the right side, can you figure those derivatives out? :D

OpenStudy (anonymous):

im not exactly sure

zepdrix (zepdrix):

Deal with each term individually :o

zepdrix (zepdrix):

For the first term, ignore the 1/2. It won't affect the differentiation process. We'll just multiply the derivative of lnx by 1/2.

OpenStudy (anonymous):

would it be 1/x

zepdrix (zepdrix):

\[\large \frac{1}{y}y'= \qquad \frac{1}{2}\cdot\frac{1}{x}+\left(\frac{x^2}{2}+7\ln\left[x^2+1\right]\right)'\]Ok good, that takes care of the first term.

OpenStudy (anonymous):

would its then be 1/2*x^2/2

zepdrix (zepdrix):

Hmmm

zepdrix (zepdrix):

For this next term you're taking the derivative of \(\large \dfrac{x^2}{2}\) Which we could write this way before taking the derivative \(\large \dfrac{1}{2}x^2\) From there, just apply the power rule.

OpenStudy (anonymous):

would it just be x

zepdrix (zepdrix):

\[\large \frac{1}{y}y'= \qquad \frac{1}{2x}+x+\left(7\ln\left[x^2+1\right]\right)'\] So the 2 came down, cancelling with the 1/2? Mm yah that sounds right.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

how would you do 7lnx^2+1

zepdrix (zepdrix):

We can pull the 7 outside, it's a constant it won't affect this process.\[\large \left(7\ln\left[x^2+1\right]\right)' \qquad=\qquad 7\left(\ln\left[x^2+1\right]\right)'\] Mmm ok this might be a little confusing at first. You'll need to get comfortable with your log derivative. You determined earlier that the derivative of \(\large \ln (x)\) is \(\large \dfrac{1}{(x)}\). The same idea follows here: the derivative of \(\large \ln\left[x^2+1\right]\) will give us \(\large \dfrac{1}{\left[x^2+1\right]}\). The entire contents of the log will go into the denominator there. But we also have to apply the chain rule, and multiply by the derivative of the inner function.\[\large \dfrac{1}{\left[x^2+1\right]}\left[x^2+1\right]' \qquad=\qquad \dfrac{1}{\left[x^2+1\right]}\left[2x\right]\]

zepdrix (zepdrix):

Oh and then of course we need that 7 from before :o\[\large 7\dfrac{1}{\left[x^2+1\right]}\left[2x\right]\]

OpenStudy (anonymous):

makes sense. is there more that needs to be done to this?

zepdrix (zepdrix):

We have a y in the denominator on the left, so we need to deal with that. Because our goal is to write our problem as y'=stuff. \[\large \frac{1}{y}y' \qquad=\qquad \frac{1}{2x}+x+7\frac{1}{x^2+1}(2x)\]Multiplying both sides by y gives us,\[\large y' \qquad=\qquad \color{royalblue}{y}\left[\frac{1}{2x}+x+7\frac{1}{x^2+1}(2x)\right]\] From here, we would like our answer to be written in terms of `x`. That means: only x's on the right side. Going back to our original equation: \[\large \color{royalblue}{y=\sqrt{x e^{x^2}}(x^2+1)^7}\] Let's plug this in for the y that's showing up in our solution.\[\large y' \qquad=\qquad \color{royalblue}{\sqrt{x e^{x^2}}(x^2+1)^7}\left[\frac{1}{2x}+x+7\frac{1}{x^2+1}(2x)\right]\]

OpenStudy (anonymous):

can this be multiplied

OpenStudy (anonymous):

if so how would it be done

zepdrix (zepdrix):

No we wouldn't want to multiply that out :) We would maybe want to simplify the part inside of the brackets though. So it's written as a single fraction.

zepdrix (zepdrix):

Er actually, if you leave it how it is, it SHOWS that you did logarithmic differentation. Because we can clearly see the derivative of each logarithm in this form. You probably don't want to simply too much further.

zepdrix (zepdrix):

This is such a longggg process when you do logarithmic differentiation. I hope I didn't make any silly mistakes in there :\

OpenStudy (anonymous):

lemme plug it in and check for any errors

OpenStudy (anonymous):

they are saying that it is incorrect

zepdrix (zepdrix):

k one sec :)

OpenStudy (anonymous):

thanks

zepdrix (zepdrix):

I'm trying to check the problem on Wolfram, but they didn't use logarithmic differentiation. So the end result looks a lot different lol.

OpenStudy (anonymous):

looking at answers similar to the problem yours seems correct. question is it xe^x^2 or is it xe^x2

zepdrix (zepdrix):

You'd probably want to format it like this, xe^(x^2)

zepdrix (zepdrix):

If we can't get it to match up for some reason, let's just put in the form that results from doing the problem the other way. I have the answer up on wolfram.

OpenStudy (anonymous):

its alright thanks for the help. i got it. i missed formatted it. i really appreciate it

zepdrix (zepdrix):

Ah ok good :) no prob

OpenStudy (cggurumanjunath):